### v1

# A review and experiences from quantum physics educators

Abstract
Students often enter physics classrooms with deeply ingrained misconceptions, typically stemming from common intuition and everyday experiences. These misconceptions present significant challenges for educators, as students are often resistant to information that conflicts with their preconceptions. As a result, traditional instructional methods often fail to address misconceptions. The first aim of this manuscript is to summarize the existing literature on misconceptions in university physics. This resource for instructors reviews misconceptions’ sources, diagnoses, and remediation strategies. Like most physics education research, the majority of this literature has concentrated on classical physics. However, quantum physics poses unique challenges because its concepts are far removed from everyday experiences and intuition. This uniqueness signals the need to ask how well the strategies developed for addressing misconceptions in classical physics apply to quantum physics. This need is underscored by the recent surge of people from diverse backgrounds entering quantum physics because of the growing significance of quantum technologies in fields such as computing, cryptography, and materials science. To help answer this question, we conducted in-depth interviews with quantum physics instructors at the University of Waterloo who have collectively taught over 100 university quantum physics courses. These interviews explored the nature of common misconceptions in quantum physics, their origins, and effective instructional techniques to address them. We highlight specific misconceptions, such as misunderstanding of entanglement and spin, and successful teaching strategies, including “misconception-trap quizzes.” We integrate insights from the literature review with our interview data to provide an overview of current best practices in addressing physics misconceptions. Furthermore, we identify key research questions that warrant further exploration, such as the efficacy of multi-tier tests in quantum physics and developing a cohesive quantum curriculum. This paper aims to inform educators and curriculum developers, offering practical recommendations and setting a research agenda to improve conceptual understanding in classical and quantum physics.
I Introduction
Physics is filled with concepts that challenge common intuition and are difficult for students to grasp. For example, many students mistakenly believe that heavier objects fall faster than lighter ones or that a constant force is needed to keep an object in motion. Such misconceptions often arise from everyday experiences and intuitive reasoning, leading to persistent barriers in learning even after formal instruction [1, 2].
Misconceptions can be more detrimental than gaps in knowledge. This is because students tend to resist new information that conflicts with their existing beliefs, resulting in cognitive dissonance and frustration [3, 4]. Therefore, effective teaching strategies that identify and rectify these misconceptions are essential for fostering a conducive learning environment.
However, physics educators face significant challenges in identifying and rectifying these misconceptions. Traditional instructional methods often fail to address the underlying intuitive beliefs that give rise to misconceptions [5, 6]. Not all traditional teaching methods teach for conceptual change [7]. Consequently, educators must employ innovative and targeted strategies to help students develop a more accurate conceptual framework. Fortunately, a sizeable body of research has been conducted on evidence-based strategies for identifying and addressing misconceptions. The first aim of this manuscript is to review the existing literature on misconceptions in university physics, focusing on their sources, diagnosis, and remediation strategies.
Most physics education is centred on classical physics; naturally, most physics education research also focuses on classical physics. This trend extends to the study of misconceptions. For instance, Gilbert and Watts [1] reviewed 33 studies from 1972 to 1982 on identifying misconceptions, and Soeharto et al.  [2] reviewed another 111 articles from 2015 to 2019. Of the over 40 misconceptions identified in these studies, none were on quantum physics, highlighting a potential gap in the research.
Quantum physics introduces a unique set of challenges due to its abstract and non-intuitive nature [8, 9, 10]. Phenomena such as wave-particle duality and entanglement are far removed from everyday experiences and defy classical logic. This uniqueness underscores the need for increased attention to identifying and remedying misconceptions in quantum physics. The importance of addressing these misconceptions is further amplified by the growing significance of quantum technologies in fields such as computing, cryptography, and materials science. This growth has resulted in learners from a broad range of academic backgrounds looking to learn the basics of quantum physics.
Strategies designed to diagnose and remedy misconceptions in classical physics often rely on analogies and experiences from the macroscopic world, which do not always translate well to quantum phenomena. Consequently, while it is essential to determine which ideas and tools from classical physics education can be adapted for quantum physics, it is equally important to recognize that some may not directly apply. To explore the applicability of these strategies, we conducted interviews with quantum physics instructors at the University of Waterloo, a leading institute for quantum physics which houses the Institute for Quantum Computing and has strong affiliations with the Perimeter Institute for Theoretical Physics. Collectively, these instructors have taught over 100 quantum physics courses. Since they are on the front lines of teaching and observing student misconceptions, they provide valuable insights into the effectiveness of existing educational tools and the unique challenges faced in teaching quantum concepts. These interviews aim to identify which educational strategies can be ported over from classical to quantum physics. This is the second aim of this manuscript.
The rest of the manuscript is organized as follows. Section II presents our review of the literature on addressing misconceptions in university physics. Instead of providing an exhaustive survey, we highlight key results and illustrative examples of each. A summary of the interviews with quantum physics instructors is presented in Section III. By integrating insights from these interviews with the existing literature on classical physics misconceptions, we aim to provide an overview of best practices for addressing misconceptions in classical and quantum physics and identify research opportunities. This integrated discussion is presented in Section IV.
II Review
This section reviews the literature on misconceptions in university physics education, beginning with a review of the broader context of potential sources of misunderstanding (Section II.1). In Section II.2, we explore the origins of these misconceptions. Next, we examine methods for diagnosing misconceptions in Section II.3. Finally, in Section II.4, we discuss strategies for remedying these misconceptions. We propose that these remediation strategies can be categorized into three main types: experiment-based, discussion-based, and reflection-based. This categorization helps to identify common elements and key factors contributing to their effectiveness.
II.1 Misconceptions among other sources of misunderstanding
The study of misconceptions begins with defining them precisely and distinguishing them from other sources of misunderstanding. In their 1983 review, Gilbert and Watts noted the initial lack of consensus on the definition of a misconception within the field [1]. We review two well-known typologies for categorizing misunderstandings and simplify them into a two-bin system: misconceptions, stemming from incorrect existing knowledge, and gaps in prior knowledge.
One prominent typology was proposed by the National Research Council  [11]. They presented five sources of misunderstandings:
  • (1a)
    Preconceived notions: Intuitive but often incorrect understandings of scientific concepts based on everyday experiences.
  • (1b)
    Non-scientific beliefs: Views learned outside the scientific community that often conflict with scientific evidence.
  • (1c)
    Conceptual misunderstandings: Errors arising from incorrectly relating scientific information, leading to unresolved paradoxes or conflicts.
  • (1d)
    Factual misconceptions: Incorrect facts learned in childhood, often perpetuated by authoritative figures like parents and teachers.
  • (1e)
    Vernacular misconceptions: Misunderstandings due to the different meanings of words in scientific versus everyday language.
Liu and Fang’s meta-analysis [12] presents another well-studied classification. Their work reviewed 60 papers on misunderstandings about force and acceleration across various educational levels, identifying 38 misunderstandings related to force and 15 related to acceleration. These were categorized into four primary causes:
  1. (2a)
    Incomplete or partial understanding: Occurs when students do not fully comprehend key concepts, leading to gaps in their knowledge.
  2. (2b)
    Preconceived misunderstandings: Misunderstandings carried over from a student’s life experiences.
  3. (2c)
    Wrong interpretations and comprehensions: When students understand a concept in isolation but are unsure how to apply it.
  4. (2d)
    Vernacular misunderstandings: Arise from students interpreting scientific terminology based on everyday language usage.
Our category “gaps in knowledge” aligns with 2a, and “misconceptions” with 1a-1d and 2b-2c. Which binning 1e or 2d goes under may depend on context, but it seems more likely to be a knowledge gap. For example, translating a word incorrectly is a vernacular misunderstanding that can likely be corrected easily without leading to deeply ingrained incorrect beliefs. Having clearly defined misconceptions and tied our definition to the existing ones used in the literature, we turn to their sources.
II.2 Sources of misconceptions
Misconceptions can arise from three sources:
  • Sources of knowledge: Misconceptions can arise from inaccurate information from sources students believe they should trust to some extent. This includes teachers themselves, textbooks, videos, and other literature.
  • Teaching Methods: Ineffective teaching strategies can contribute to forming and perpetuating misconceptions. This includes complicated entry points into a subject or poorly designed teaching activities.
  • Students: Even in the presence of accurate sources of knowledge with sound teaching methods, students can develop misconceptions. This can be from their prior experiences, incorrect reasoning, stage of development, abilities, or interest in the subject.
Our three-bin system aligns with Ref. [13]’s if one group “Textbooks and Literature” with “Teachers.” We do not distinguish these because the line between them is inherently unclear. Textbooks and literature were written by teachers, and teachers will provide their lecture notes to students. Furthermore, students often rely on online teachers, such as Khan Academy, to fill gaps. Which category would this fall into? Our system avoids this ambiguity.
When students do not yet understand a subject, they cannot distinguish correct from faulty sources of knowledge. Kulgemeyer and Wittwer examined the emergence of misconceptions during self-directed learning, particularly through the use of physics explainer videos, which have become increasingly popular over the past two decades [14]. Their study assessed the influence of these videos on students’ understanding by dividing 149 physics learners into two groups. One group viewed a video containing misconceptions about the concept of force, while the control group watched a scientifically accurate video. Both videos were similar in terms of comprehensibility and duration, differing primarily in content accuracy. The results revealed that the group exposed to misconceptions developed a comparable level of perceived understanding to the control group but acquired more misconceptions and less scientific knowledge. This finding highlights the critical importance of evaluating the accuracy of educational content in textbooks and other learning resources to prevent the dissemination of misinformation.
Teaching methods that are often effective in filling gaps in knowledge can, in fact, lead to misconceptions when not implemented conscientiously. For example, Cook discussed visual representations’ role in this process [15]. According to Cook, when visual aids are presented without adequate prior knowledge, they can lead to misunderstandings about scientific concepts. The risk of misconceptions increases when the visuals overwhelm the learner’s cognitive capacity. Since prior knowledge is essential for reducing cognitive load, its absence can lead individuals with limited prior knowledge to focus on superficial features of the representations and derive incorrect conclusions. The careful design of visual materials and the use of multiple representations help mitigate this problem. While Cook’s perspective is compelling, it warrants further experimental investigation. Looking ahead, our interviews revealed many instances where this phenomenon occurred in the context of quantum physics. This is potentially due to the abstract nature of quantum concepts exacerbating the impact of vague instructional materials.
II.3 Diagnostic Methods
Being mindful of sources of misconceptions will reduce the number of students’ misconceptions, but some will emerge. To help students address these misconceptions, we need tools to diagnose them. The commonly used tools are interviews, open-ended questions, certainty of response index (CRI) tests, and multiple-tier tests. The first two strategies are typically used in courses with few students, and the latter two in courses with many students.
Interviews and open-ended questions
Interviews involve one-on-one interactions between an interviewer and a student. These sessions aim to uncover the student’s reasoning on specific topics, thereby revealing underlying misconceptions. The primary strengths of interviews include their ability to provide in-depth insights into students’ cognitive processes, their adaptability in real-time questioning, and their potential to generate rich, qualitative data. These advantages enable a nuanced assessment of students’ conceptual understanding.
However, interviews also present several limitations. They are time-intensive, both in execution and subsequent analysis, which restricts their scalability for large student cohorts. Additionally, the quality of data collected can be significantly influenced by the interviewer’s skill and potential biases, introducing subjectivity into the assessment. The analysis process itself can further contribute to this subjectivity. Moreover, a certain level of trust between the student and the interviewer is necessary to ensure candid responses.
With Open-ended questions, students articulate their understanding and reasoning freely without the constraints of predefined answer choices. This format encourages comprehensive responses, offering deeper insights into students’ thought processes than multiple-choice questions. Nonetheless, they share many of the same strengths and limitations as interviews. They provide depth, but analyzing open-ended responses is time-consuming, particularly in large classes. Considerable effort is required to interpret and categorize free-text answers, further introducing subjectivity into the evaluation. Additionally, students may not provide detailed answers unless incentivized, which can limit the effectiveness of this method compared to oral interviews.
CRI and multiple-tier tests
The foundational concept of "multiple-tier" physics tests was proposed by Hasan et al. [16]. They adopted the Certainty of Response Index (CRI), originally utilized in social sciences, to assess students’ confidence in their multiple-choice answers on a scale from 0 (total guess) to 5 (complete confidence). Incorrect answers with low CRI scores indicate knowledge gaps, while incorrect answers with high CRI scores suggest the presence of misconceptions. Hasan et al. applied this approach first in an undergraduate classical mechanics course. They found that average CRI values provided insights into the overall understanding of the class, guiding instructional strategies. This approach, also called a “two-tier multiple-choice” test, has become a cornerstone in diagnosing misconceptions.
However, the CRI method has its limitations. For example, students might confidently choose the correct answer for incorrect reasons. To address this issue, the four-tier diagnostic test [17, 18] was developed. This method requires students to select a multiple-choice answer (first tier) and rate their confidence (second tier), then provide a justification for their choice (third tier) and rate their confidence in that justification (fourth tier). Although more time-consuming, this comprehensive approach better identifies misconceptions by evaluating the answers and the reasoning behind them. Additionally, "three-tier multiple-choice" tests, which add only the third tier, have been experimented with to balance thoroughness and practicality. Another limitation of the CRI is that students may understand the concepts but lack confidence in their answers. Researchers have suggested recognizing a separate category of students who grasp the material but are uncertain about their understanding [19, 20]. Identifying this population is necessary for developing tools to support them.
Multi-tier tests have eclipsed more long-form evaluations in use. To illustrate this, we reference three review articles [21, 2, 22]. Gurel et al. [21] analyzed 273 publications from 1980 to 2014, finding that interviews were used in 53% of studies, open-ended questions in 34%, and CRI and multiple-tier tests in the remainder. A subsequent review by Soeharto et al. [2] covering 111 studies from 2015 to 2019 showed a shift: interviews in 11%, open-ended questions in 24%, and the rest using CRI and multiple-tier tests. Resbiantoro et al. [22] reviewed 72 articles from 2005 to 2020, confirming this trend with interviews in about 10%, open-ended questions in 28%, and the rest using CRI and multiple-tier tests. These reviews highlight transitioning from interactive methodologies to more standardized testing formats in assessing misconceptions. The exact reason for this is unclear. Multi-tier tests may be more effective, or physics education research may have drifted towards studying classrooms with more students.
II.4 Remediation Strategies
After a misconception is identified, it can be addressed. This requires effective strategies grounded in educational research to foster conceptual change [7]. In this section, we propose a typology that categorizes these strategies into three groups: reflection-based, discussion-based, and experiment-based approaches. These categories highlight the underlying principles shared amongst various strategies. Examples for each category are summarized in Table 1.
Strategy Explanation
Reflection-based
Concept Mapping Learners create maps linking concepts by connecting words or phrases, illustrating relationships between ideas. This helps organize and structure knowledge, facilitating better comprehension and recall.
Journaling Learners record their thoughts, understanding, and progress over time, enhancing reflection and deepening learning.
Predict-Observe-Explain Learners make predictions about an outcome, observe an experiment to see what actually happens, and then explain the results. This helps clarify and modify their understanding based on observed evidence.
Conceptual-Change Text Learners are presented with text that challenges existing beliefs and encourages cognitive conflict. These texts help reshape understanding toward scientifically sound concepts.
Discussion-based
Group Discussions Learners participate in structured or informal discussions among learners or instructors to explore ideas, clarify understandings, and negotiate interpretations. These discussions promote critical thinking and deeper comprehension through collaborative dialogue.
Listening Activities Learners listen to others discuss concepts. These can be pre-recorded conversations or live ones in class. Hearing the conflict of other ideas helps students in refining their own misconceptions.
Experiment-based
Laboratory Learners engage in hands-on experiments in a laboratory setting, enabling them to engage with processes relevant to the subject matter directly. This method supports experiential learning and helps students connect theory with practical application.
Simulations Learners use computer simulations to mimic real-world processes or experiments, allowing them to interact with complex systems in a controlled, virtual environment. This helps in understanding abstract concepts and testing hypotheses without the constraints of physical experiments.
Table 1: Summary of educational strategies categorized into reflection-based, discussion-based, and experiment-based approaches.
Reflection-based strategies
Reflection-based strategies encourage students to introspect and critically reassess their knowledge, integrating new and accurate information. These methods, often involving individual activities, can also include collaborative elements. Examples of reflection-based strategies include concept mapping [23], journaling [24], model-based teaching [25], predict-observe-explain activities [26], and conceptual change text [27, 28, 29, 30].
Hein’s 1999 study highlighted how reflection-based strategies with writing components can effectively identify and remedy misconceptions [24]. They introduced a “folder activity” in an algebra-based introductory physics course for non-science majors at American University in Washington, DC. This course often enrolls students with limited mathematical skills and no prior physics education. The folder activity, conducted 5–10 times per semester, includes prompts for short-answer questions related to course concepts. Students use their class notes to formulate responses, promoting a deeper understanding through the process of writing. Additionally, some prompts require students to create multiple-choice questions and justify their answers. This approach not only helps students clarify their own understanding but also allows instructors to identify common misconceptions for further discussion. Hein’s argues that this method is effective both as a diagnostic tool and as a strategy for correcting misconceptions.
Reflection activities do not require long-form writing to be effective; even trying to reflect helps. In a study by Miller et al.,  [26], the impact of the predict-observe-explain method on student learning in introductory physics courses was examined. They focused on the subject of mechanics and of electricity and magnetism across two universities. Students were asked to predict the outcomes of 22 different demonstrations before observing them. Researchers recorded the students’ predictions and observations, analyzing the data immediately after the demonstrations and at the end of the semester. They found that about 20% of the observations did not match the actual outcomes due to a misconception. However, students who made predictions were about 20% more likely to correctly perceive the demonstration results, regardless of whether their predictions were accurate. Their results suggest that the act of making predictions enhances students’ observational accuracy and helps them retain the correct outcomes.
Discussion-based strategies
Discussion-based activities use social interaction to boost learning. Students encounter various perspectives and reasoning methods by engaging in discussions and debates with peers and instructors. These strategies have been effective in multiple formats, including in-person interactions [31], online discussions [32], and even through listening to recorded conversations [33].
Informal discussions are often employed in classes and can be highly effective. However, these activities can benefit from higher levels of structure. Leinonen et al. [31] work demonstrates a more structured example. They studied the effect of a discussion-based strategy in an introductory thermal physics course at the University of Eastern Finland. It was a one-hour session divided into three phases: individual work, hinting, and peer discussion. The effectiveness of the intervention was evaluated by assessing students’ conceptual understanding before, during, and after the session. This assessment included written explanations from all the students and recordings of peer discussions from five pairs of volunteers. The study focused on 65 students out of the 100-120 enrolled in the course, covering chemistry, mathematics, physics, and computer science majors. Initially, 75 students participated in the pretesting, but the analysis centred on the 65 who completed all stages of the intervention, ensuring consistency across the pretesting, intervention, and post-testing phases. The intervention significantly improved students’ understanding of thermal physics concepts, especially after the peer discussion phase. For instance, the percentage of correct answers to a question about work in an isobaric process increased from 52% in the individual phase to 80% after peer discussion. Similarly, following peer discussions, correct responses for the net work done in a cyclic process rose from 28% to 51%, and for heat in a cyclic process, from 15% to 40%.
Experiment-based strategies
The final category, experiment-based activities, emphasizes “hands-on” engagement, allowing students to interact directly with physical or simulated systems. These tangible experiences help students visualize and understand complex concepts, receive live feedback, and reconstruct their knowledge based on observed outcomes. Experiment-based learning includes laboratory experiments [34] and simulation-based experiments, which have gained popularity since the early 2000s. These simulations include interactive tools [35, 36, 37, 38, 39, 40] and virtual reality technology [41].
The simulation-based experiments provide the opportunity to bake additional teaching principles into the experiment. This is exemplified nicely by the Virtual Physics System (ViPS) developed by Myneni et al. [35] ViPS is a software designed to enhance students’ understanding of pulleys. ViPS integrates simulation and tutoring to address common misconceptions. The ViPS process includes three stages:
  1. 1.
    Pre-Test Phase: Students work on problems that reveal their misconceptions about pulleys. This provides ViPS with a baseline understanding.
  2. 2.
    Tutoring Phase: ViPS customizes the experience to address identified misconceptions. Students go through guided problem-solving, receiving real-time feedback and hints, with guidance levels adjusted according to their performance. This is done to adhere to the zone of proximal development principle [42].
  3. 3.
    Post-Test Phase: Students’ comprehension is then assessed post-instruction to verify the rectification of misconceptions.
ViPS uses a dynamic student model to track and adapt to each learner’s progress, ensuring personalized feedback and challenges. This adaptability tailors the educational experience to the student’s evolving understanding. This capability goes beyond what is typically possible in laboratory-based experiments.
Myneni et al. compared ViPS with traditional laboratory experiments. The study involved 12 engineering majors and 210 preservice elementary teachers, divided into groups experiencing various combinations of virtual and physical pulley interactions. Virtual and physical sequences improved understanding, but virtual interactions alone were more impactful than physical ones alone, emphasizing the potential of simulation-based learning tools in educational settings.
III Experiences from quantum physics educators
Most of the tools for identifying and addressing misconceptions were developed for classical physics. Since quantum physics education presents unique challenges, we need to consider the applicability of these tools to this subject. We conduct such an evaluation through long-form interviews with quantum physics instructors at the University of Waterloo. This study aims to identify misconceptions in quantum physics and assess effective tools for diagnosing and rectifying these misconceptions.
III.1 Methodology Overview
First, we outline our methodology for identifying instructors and conducting interviews. The University of Waterloo, renowned for its quantum physics research and extensive course offerings, was chosen for this study. It houses the Institute for Quantum Computing and is closely affiliated with the nearby Perimeter Institute for Theoretical Physics. A testament to the university’s focus on quantum physics is the volume of courses it offers on the subject. Listed in Appendix A, the university offers 16 undergraduate and 37 graduate courses on solely quantum physics and 9 undergraduate courses covering quantum physics.
We identified faculty members who taught the 16 undergraduate courses via online course listings and departmental contacts. Emails were sent to these faculty members requesting interviews, resulting in nine agreements. Through these nine, another faculty member was identified to give an interview about their experience teaching one of the nine undergraduate courses partially covering quantum physics. Additional interviews were secured with two individuals from the Institute for Quantum Computing and Perimeter Institute with years of experience teaching quantum physics to undergraduate students. The twelve instructors have collectively taught over 100 university quantum physics courses and many summer schools, reaching thousands of students.
With no detailed guide for conducting physics education research interviews, we relied on guides from other fields [43, 44, 45, 46]. The interviews aimed to identify misconceptions in quantum physics and effective tools for addressing them. Five core questions were developed, with flexibility for follow-up questions based on responses. Verbal consent for recording was obtained, and confidentiality and withdrawal rights were explained to each interviewee. Interviews were scheduled for 30 minutes each, conducted via Zoom, and many were extended beyond the scheduled time. The insights gathered from these interviews are synthesized and presented in a cohesive narrative in the following subsection.
III.2 Synthesis of Insights
We present each question and list the various answers instructors gave. With few exceptions, each answer is an amalgamation of comments given by multiple instructors.
What are university students’ common misconceptions about quantum physics?
Entanglement.—Students often believe that entanglement implies faster-than-light communication. They think a local change in system $A$ will instantaneously cause a change in another system entangled with $A$. This misconception is frequently attributed to portrayals in popular educational materials, which often dramatize quantum phenomena without accurate scientific context. For example, students might assume that measuring one particle’s spin immediately determines its entangled partner’s spin, overlooking the fact that no information is transmitted in this process.
Spin.—Students frequently misunderstand spin as arising from an object physically spinning. They struggle to grasp what it means for spin to be an intrinsic property of particles, unrelated to any literal rotation. This misconception is often reinforced by incorrect depictions of electrons as rotating charged spheres in educational materials. For example, students may visualize an electron’s spin as similar to a spinning top.
Observer effect.—A common belief among students is that a conscious observer is necessary for the collapse of the wave function. This misconception is perpetuated by textbook illustrations showing an eyeball next to a quantum system and by sensationalized portrayals in popular media. For example, students might ignore the role of measurement devices in determining a particle’s state if there is no human observation.
Uncertainty in a system’s state.—Students tend to erroneously believe that the state of a quantum system is always uncertain. For example, students may assume the energy of the Hamiltonian’s eigenstates is uncertain. Confusion about the Heisenberg Uncertainty Principle may be tied to this. Some students interpret the uncertainty principle as implying that quantum systems are always fundamentally unpredictable rather than understanding it as a limit on the precision of simultaneous measurements.
Density matrices.—Students often equate classical uncertainty in a system’s state with quantum uncertainty. A lack of clear distinction between classical and quantum probabilities in educational resources contributes to this misunderstanding. Misunderstanding the nature of mixed states versus pure states can lead students to confuse the probabilistic interpretations in classical and quantum contexts.
Quantization.—Students mistakenly think that quantization is unique to quantum mechanics, overlooking the fact that classical systems can also be quantized. Classical systems, such as energy levels in a harmonic oscillator, can exhibit quantized properties.
Wave-particle duality.—Students struggle to reconcile the dual nature of particles without conflating the concepts, misinterpreting the principle as literal simultaneity rather than a complementary description. Ambiguous language in textbooks often leads to this confusion.
Quantum theory of light.—Students often confuse the energies of different modes with the number of photons in a mode during exercises. This confusion may be exacerbated by the use of similar ladder-like diagrams to represent both concepts. Misinterpreting photon interactions and energy levels can lead to significant misunderstandings in quantum optics exercises.
What are the sources of these misconceptions?
The instructors collectively identified eight sources of misconceptions. The first four pertain to teaching methods, the next three to educational materials, and the last to the students.
Traditional introductions to quantum mechanics.—Misconceptions can stem from traditional introductions to quantum mechanics. One traditional approach is starting with wave functions. Multiple instructors pointed out that doing so instead of working with Dirac notation and discrete systems increases the mathematical barrier to entry in quantum mechanics. More instructors are now adopting the Dirac notation approach, which has been beneficial in reducing initial learning difficulties.
Another problematic entry point is starting with the Stern–Gerlach experiment. This method, used in some standard quantum physics texts, introduces many complex concepts simultaneously, such as spin and measurement collapse. These concepts are common sources of misconceptions. Thus, starting with the Stern–Gerlach experiment may necessarily overload a student as they are first learning the subject.
Distinguishing experimental observations from interpretations.—Another source of misconceptions arises when instructors do not clearly distinguish between what is experimentally observed in quantum physics and how these observations are interpreted. Students are naturally curious about the fundamental behaviour of the universe and seek explanations, such as why measurements project a system onto an eigenstate. However, quantum physics offers multiple interpretations, and instructors can only state the experimental predictions with certainty. Mixing interpretations with experimental facts can lead to misconceptions, such as the belief that a conscious observer is necessary to collapse the wave function. While discussing interpretations can be engaging and beneficial for students, it is crucial to maintain a clear distinction between ontological facts and interpretative frameworks to prevent misunderstandings.
Prolonged disconnect from physical systems.—Quantum mechanics inherently requires mathematical proficiency, and early in one’s studies, much of the work involves rigorous mathematical manipulation. However, a prolonged focus on mathematical abstraction as one progresses to upper-year courses can lead to shallow understandings of physical systems and misconceptions. For example, students often fail to recognize the physical constraints that should be applied to their mathematical work. To demonstrate that the flux in a ring is quantized, students need the mathematical framework of Stoke’s theorem and physical arguments related to boundary conditions. Additionally, students frequently lack awareness of the approximate scales at which different quantum effects become relevant. For instance, when given a Hamiltonian, they might not know whether a perturbation will break the degeneracy of the Hamiltonian, or if the distance between ions in a trap will cause interference.
Jargon.—One instructor who teaches students from various departments highlighted the substantial amount of technical language or jargon in quantum physics that can be challenging to students from other fields, such as engineering or mathematics. This linguistic barrier can also contribute to misconceptions. Simplifying language and providing clear explanations of jargon can help bridge this gap and reduce misunderstandings.
Popular science media.—The mysterious allure of quantum physics has inspired numerous popular science media. This media, along with other educational materials, often dramatize or oversimplify quantum phenomena, leading to widespread misconceptions. For instance, the belief that entanglement enables faster-than-light communication was most often attributed to popular media. Educators must clarify and debunk these popular misconceptions in the classroom to ensure students understand accurately.
Ambiguously worded textbooks.—Beyond popular media, ambiguously worded textbooks were also identified as a source of confusion. For example, the concept of wave-particle duality is often misrepresented in textbooks, which might state that a photon is both a wave and a particle, whereas the reality is more nuanced. Using precise language and providing nuanced explanations can help mitigate this issue.
Ambigous figures.—Various ambiguous figures were mentioned by instructors. These include images of something rotating to depict spin, an eyeball beside a system to depict projective measurements, and rope-like drawings connecting systems to depict entanglement.
Inappropriate use of problem-solving tools.—One instructor provided an example of a misconception stemming from student practices. This involved students using a familiar problem-solving toolbox inappropriately in more complex settings, thereby misunderstanding the limitations of the method. However, it was acknowledged that this issue might originate from inadequate instruction by a previous teacher. Teaching the limitations of different problem-solving tools and methods can help students apply these tools appropriately in various contexts.
What are your strategies for identifying misconceptions?
Identifying trends in formative assessment errors.—Instructors commonly analyze student responses to exercises to detect misconceptions. By examining specific types of errors or patterns in student answers, they can identify areas where students consistently struggle.
Informal discussions with subsets of students.—Informal conversations during class or office hours with students are another key strategy. Although involving only a subset of students, these discussions provide insights into common misconceptions. Instructors noted that even these limited interactions help them identify issues that can be addressed with the entire class. The spontaneous nature of these discussions often reveals student thinking processes and misunderstandings that may not surface in formal assessments.
Compiling common misconceptions.—One instructor highlighted the value of maintaining a comprehensive list of student misconceptions. This list is compiled from both personal teaching experience and shared insights from colleagues. The instructor found that misconceptions remained consistent from year to year, reducing the need for continual reassessment. This consistency allows the instructor to anticipate and proactively address common misunderstandings in future courses.
What are your strategies for addressing misconceptions?
In-class quizzes with “misconception traps”.—Several instructors utilize quizzes designed to address misconceptions. These quizzes often include "trap answers" that highlight common misunderstandings. By polling the class using electronic clickers or polling apps, instructors can reveal the range of answers and the prevalence of certain misconceptions in real-time. For instance, one instructor presents various scenarios of measurements being done on a quantum system to test students on whether they think a conscious being needs to observe an experiment for the wave function to collapse.
After polling, instructors display the results on a screen, showcasing the distribution of answers. This visual representation helps students see they are not alone in their misconceptions. A particularly effective technique involves asking students to justify an answer they did not choose. This process encourages students to engage critically with the material and understand different perspectives. It also makes students more comfortable speaking since they are not necessarily justifying their own answers.
Following this, students are given time to discuss their reasoning among themselves before being re-polled. This peer instruction phase allows students to articulate their thoughts and learn from their classmates. The instructor then revisits the question and polls the class again. This method almost always results in a consensus on the correct answer, demonstrating the power of peer instruction and collaborative learning in addressing misconceptions.
Class discussions with conceptual-change text.—Another effective strategy is employing discussions integrated with conceptual-change text. In this approach, students are presented with a scenario prone to misconceptions and asked to predict the outcome. For example, one instructor presented an interferometry experiment and asked students to predict what changes when, for example, one path is blocked or the phase of one phase is changed. Erroneous conceptions are revealed and corrected through dynamic, back-and-forth dialogue between the instructor and the class. One instructor noted, "I get them to the point where they don’t know what to believe anymore. Then, when they’re at that point, they’re ready to let go of their misconception." This strategy underscores the importance of first identifying the misconception, as it helps students remember and learn from their errors.
Video demonstrations or live experiments.—Three instructors found that video demonstrations or, when possible, live experiments effectively address misconceptions. These activities can also take the form of predict-observe-explain exercises or be accompanied by straightforward explanations. For example, an instructor might show a laser going through slits to show the wave-like interference patterns. Demonstrations also evoke a sense of awe and engagement in students, enhancing their overall learning experience. However, most instructors pointed out the limited options for this in quantum physics classrooms.
Besides misconceptions and gaps in knowledge, what else causes students to make mistakes?
Three instructors said students make mistakes for reasons outside of misconceptions and gaps in knowledge. All three pointed to deficiencies in prior knowledge.
Insufficient mathematical background.—Many second-year students lack the mathematical training required for quantum mechanics. The subject demands an understanding of complex numbers, linear algebra, and a transition to abstract Hilbert spaces. Without this strong mathematical foundation, students struggle to develop the intuition to correctly verify their answers and apply quantum concepts.
Insufficient quantum physics background.—Courses that include significant quantum mechanics content often admit students without sufficient background knowledge. This issue is particularly prevalent in courses that do not exclusively teach quantum physics, such as statistical mechanics courses. One such course at Waterloo, for example, is almost evenly divided between teaching classical and quantum statistical mechanics. The scheduling of these courses sometimes fails to account for the necessary preparatory coursework, leading to students encountering complex quantum topics without a solid foundational understanding.
Other miscellanenous comments
Several insightful side conversations emerged from the interviews, providing valuable perspectives on addressing misconceptions.
Types of misconceptions.—Instructors identified two types of misconceptions: unintended and detrimental misconceptions and what can be termed “necessary and temporary misconceptions.” Unintended and detrimental misconceptions are incorrect beliefs that harm student understanding and have been the primary focus of this manuscript. In contrast, necessary and temporary misconceptions are simplifications that, while technically incorrect, are pedagogically useful. A classic example is the progression of atomic models taught from dense billiard balls to the "plum pudding" model and finally to orbitals and electron clouds. This illustrates how these misconceptions can serve as beneficial stepping stones in learning. These simplifications help students gradually build their understanding, making complex concepts more accessible.
Bad misconceptions may still be helpful.—Many conversations led to whether misconceptions can be more useful than harmful. Misconceptions, while adding inertia to learning, can provide a foundation upon which to build. Some argued that having an initial, albeit incorrect, conception allows for a more effective learning process where instructors can explain why a misconception is wrong and why a more accurate understanding is better. This contrasts with the challenge of teaching students with no prior conception.
Specialization in quantum physics:.—The importance of tailoring quantum mechanics education to the audience and its goals was emphasized. Mathematics students may approach quantum mechanics primarily through the lens of linear algebra with additional rules. These students could go on to prove useful results about, for example, quantum algorithms without needing knowledge of, say, scattering. Quantum physics research requires a diverse set of skills, and specialization is necessary. Whether this level of specialization is relevant at the undergraduate level is unclear. However, the question is then raised on what to include in undergraduate curriculums as the number of quantum physics education courses expands.
III.3 Limitations of the study
In this work, we focused on one institute that is a hot spot for quantum physics education. This intense concentration of experiences made Waterloo a prime location for analyzing quantum physics education. However, Waterloo is one of Canada’s most competitive physics programs and is recognized as a global leader in quantum physics. As such, the student population is likely biased to have a stronger physics background than the average student. Furthermore, the instructors likely understand the content more deeply than those at a randomly chosen university. Future research should aim to include a larger and more diverse sample of instructors from various institutions and cultural backgrounds to enhance the generalizability of the findings.
The study relies heavily on qualitative data from interviews, which, while rich in detail, may introduce subjectivity and bias. The interpretations and conclusions drawn are influenced by the educators’ and interviewer’s perspectives and experiences. Including quantitative measures, such as surveys or pre- and post-assessments of student performance, could provide a more balanced view and validate the effectiveness of the suggested strategies.
IV Discussion
Physics educators often face challenges due to ingrained student misconceptions. Classical and quantum physics present unique hurdles, as highlighted by the literature and reinforced by our interviews with quantum physics educators at the University of Waterloo. This final section synthesizes many of the conceptual insights from our review and discussions and identifies key research questions. However, one should refer to the main text for a detailed list of strategies.
IV.1 Integration of insights from literature and interviews
Instructors can save considerable time by familiarizing themselves with the common misconceptions in their courses that have already been identified in prior studies. This knowledge allows instructors to shift where and how they emphasize topics in class. Many studies listing common misconceptions exist for classical physics, and this study presents numerous examples in quantum physics.
Instructors can help limit misconceptions before they ever enter a student’s framework. To do so, instructors can review educational materials to check for accuracy before assigning them and critically analyze their teaching methods regularly. Our interviews with educators highlighted weaknesses in various quantum physics teaching strategies. However, many of these strategies have existed for decades, and their issues have only been identified through critical analysis.
Despite an instructor’s best efforts, misconceptions are inevitable. To rectify them, we need methods to identify them. The literature contains a spectrum of approaches. On one end are interviews and open-ended questions, which provide considerable depth of student understanding but are costly in time. These are typically best for smaller classes but, if necessary, can be adapted for large courses. This can be done with some form of random sampling or by hiring additional teaching assistants. On the other end of the spectrum are multi-tier tests, which are efficient and effective but come at the cost of depth. Interestingly, the quantum educators we spoke to rely more on time-consuming approaches and employ the mentioned strategies to mitigate the time demand.
The literature suggests several effective strategies for addressing misconceptions we categorized with self-explanatory names as reflection-based, discussion-based, and experiment-based (see Table 1). The main text includes detailed examples of each. We note that experiment-based strategies are less used when teaching quantum physics. However, this is due to their lack of availability, not lack of efficacy.
IV.2 Key research questions
Multi-tier tests have become prevalent for identifying misconceptions in physics education but are less commonly used in quantum physics education. Investigating their effectiveness in this field is a promising research opportunity. Furthermore, developing a pool of multi-tier questions based on identified misconceptions could also be valuable for the quantum education community.
The quantum physics misconceptions identified by educators at Waterloo are largely conceptual, contrasting with the primarily mathematical and technical misconceptions identified by Styer in 1996 [47]. A systematic study comparing the relative importance of addressing these different types of misconceptions may also prove helpful in identifying effective remediation strategies.
Instructors highlighted that deficiencies in prior knowledge, particularly in mathematical training and foundational quantum physics, significantly impact student performance. The rapid development of quantum physics education seems to have outpaced curriculum design and course scheduling. Aligning mathematical and physics curricula to ensure continuity and coherence is crucial. Developing cohesive quantum physics education programs is another area that merits exploration, with studies like Asfaw et al. [48] providing valuable frameworks. Related to this point, the tailoring of quantum physics education to different student audiences’ specific needs and goals is vital. This could be done through adaptable curricula that provide a solid foundation while allowing specialization to prepare students for diverse career paths in quantum physics and related fields.
Simulations have proven incredibly effective in teaching classical physics concepts and may be equally valuable in quantum physics education. Simulating phenomena such as the double-slit experiment or entanglement can give students a more intuitive understanding of these abstract concepts. Developing simulated experiments for quantum classrooms is another impactful area of research. For example, resources like IBM’s Qiskit and Xanadu’s Pennylane for simulating quantum algorithms exist and could be valuable tools for quantum computing education.
Acknowledgements.
This article is part of my work to fulfill the requirements for the University of Waterloo’s Centre for Teaching Excellence’s Certificate in University Teaching. This work is distinct from my PhD research [49, 50, 51, 52, 53, 54, 55, 56, 57, 58] which was funded by the Vanier CGS. However, as this article was written during this time, I would like to acknowledge and express gratitude for the support provided by the Vanier CGS. I would like to extend my thanks to Bindiya Arora, Raffi Budakian, Richard Cleve, John Donohue, Alan Jamison, Brenda Lee, Kazi Rajibul Islam, Raymond Laflamme, Eduardo Martin-Martinez, Michael Reimer, and Christopher Wilson. Your insights were illuminating, and your commitment to teaching excellence is inspiring. I would also like to thank Svitlana Taraban for her mentorship throughout the writing process and for her valuable feedback on the manuscript.
Appendix A Quantum physics courses at the University of Waterloo
Undergraduate courses on quantum physics:
  1. 1.
    AMATH 373 Quantum Theory 1
  2. 2.
    AMATH 474 Quantum Theory 3: Quantum Information and Foundations
  3. 3.
    CHEM 356 Introductory Quantum Mechanics
  4. 4.
    ECE 405 Introduction to Quantum Mechanics
  5. 5.
    NE 332 Quantum Mechanics
  6. 6.
    PHIL 252 Quantum Mechanics for Everyone
  7. 7.
    PHYS 233 Introduction to Quantum Mechanics
  8. 8.
    PHYS 234 Quantum Physics 1
  9. 9.
    PHYS 334 Quantum Physics 2
  10. 10.
    PHYS 434 Quantum Physics 3
  11. 11.
    PHYS 454/AMATH 473 Quantum Theory 2
  12. 12.
    PHYS 467/CS 467/CO 481 Introduction to Quantum Information Processing
  13. 13.
    PHYS 468 Introduction to the Implementation of Quantum Information Processing
  14. 14.
    PHYS 484 Quantum Theory 3: Quantum Information and Foundations
  15. 15.
    PMATH 343 Introduction to the Mathematics of Quantum Information
Undergraduate courses covering quantum physics:
  1. 1.
    CHEM 209 Introductory Spectroscopy and Structure
  2. 2.
    CHEM 350L Physical Chemistry Laboratory 2
  3. 3.
    ECE 457C Reinforcement Learning
  4. 4.
    PHIL 459 Studies in the Philosophy of Physics
  5. 5.
    PHYS 124 Modern Physics
  6. 6.
    PHYS 349 Advanced Computational Physics
  7. 7.
    PHYS 359 Statistical Mechanics
  8. 8.
    PHYS 435 Current Topics in Condensed Matter Physics
  9. 9.
    PHYS 444 Introduction to Particle Physics
Regularly offered graduate courses:
  1. 1.
    AMATH 673 Quantum Theory 2
  2. 2.
    AMATH 674 Quantum Theory 3: Quantum Information and Foundations
  3. 3.
    AMATH 876/QIC 845 Open Quantum Systems
  4. 4.
    AMATH 877 Foundations of Quantum Theory
  5. 5.
    CHEM 746 Quantum Chemistry
  6. 6.
    CO 781 Topics in Quantum Information
  7. 7.
    CS 766 Theory of Quantum Information
  8. 8.
    CS 867 Advanced Topics in Quantum Computing
  9. 9.
    ECE 677/QIC 885 Quantum Electronics and Photonics
  10. 10.
    NANO 707 From Atoms to Crystals, Quantum Wells, Wires and Dots
  11. 11.
    PHYS 601 Perimeter Scholars International Quantum Field Theory 1
  12. 12.
    PHYS 603 Perimeter Scholars International Quantum Field Theory 2
  13. 13.
    PHYS 605 Perimeter Scholars International Quantum Theory
  14. 14.
    PHYS 635 Perimeter Scholars International Quantum Information Review
  15. 15.
    PHYS 638 Perimeter Scholars International Quantum Gravity
  16. 16.
    PHYS 639 Perimeter Scholars International Foundations of Quantum Theory
  17. 17.
    PHYS 641 Perimeter Scholars International Explorations in Quantum Information
  18. 18.
    PHYS 644 Perimeter Scholars International Explorations in Quantum Gravity
  19. 19.
    PHYS 645 Perimeter Scholars International Explorations in Foundations of Quantum Theory
  20. 20.
    PHYS 701 Quantum Mechanics 1
  21. 21.
    PHYS 702 Quantum Mechanics 2
  22. 22.
    PHYS 703/AMATH 873 Introduction to Quantum Field Theory
  23. 23.
    PHYS 739 Quantum Many Body Physics
  24. 24.
    PHYS 760/QIC 860 Laboratory on Control of Quantum Technology
  25. 25.
    PHYS 761/QIC 861 Laboratory on Photonic Quantum Technology
  26. 26.
    PHYS 762/QIC 862 Laboratory on Low Temperature Quantum Technology and Nanofabrication
  27. 27.
    PHYS 763 Independent Project in Quantum Technology
  28. 28.
    PHYS 768 Special Topics in Quantum Information Processing
  29. 29.
    PHYS 769 Special Topics in quantum Information Processing
  30. 30.
    PHYS 785/AMATH 872 Introduction to Quantum Field Theory for Cosmology
  31. 31.
    QIC 710/PMATH 871/PHYS 767/CS 768/CO 681/AMATH 871 Quantum Information Processing
  32. 32.
    QIC 750/ECE 676 Quantum Information Processing Devices
  33. 33.
    QIC 820 Theory of Quantum Information
  34. 34.
    QIC 823 Quantum Algorithms
  35. 35.
    QIC 863 Independent Project in Quantum Technology
  36. 36.
    QIC 880 Nanoelectronics for Quantum Information Processing
  37. 37.
    QIC 890 Topics in Quantum Information

### v2

@@
-A review and experiences from quantum physics educators
+Experiences from quantum physics educators
 Abstract
-Students often enter physics classrooms with deeply ingrained misconceptions, typically stemming from common intuition and everyday experiences.
-These misconceptions present significant challenges for educators, as students are often resistant to information that conflicts with their preconceptions.
-As a result, traditional instructional methods often fail to address misconceptions.
-The first aim of this manuscript is to summarize the existing literature on misconceptions in university physics.
-This resource for instructors reviews misconceptions’ sources, diagnoses, and remediation strategies.
-Like most physics education research, the majority of this literature has concentrated on classical physics.
-However, quantum physics poses unique challenges because its concepts are far removed from everyday experiences and intuition.
-This uniqueness signals the need to ask how well the strategies developed for addressing misconceptions in classical physics apply to quantum physics.
-This need is underscored by the recent surge of people from diverse backgrounds entering quantum physics because of the growing significance of quantum technologies in fields such as computing, cryptography, and materials science.
-To help answer this question, we conducted in-depth interviews with quantum physics instructors at the University of Waterloo who have collectively taught over 100 university quantum physics courses.
-These interviews explored the nature of common misconceptions in quantum physics, their origins, and effective instructional techniques to address them.
-We highlight specific misconceptions, such as misunderstanding of entanglement and spin, and successful teaching strategies, including “misconception-trap quizzes.” We integrate insights from the literature review with our interview data to provide an overview of current best practices in addressing physics misconceptions.
-Furthermore, we identify key research questions that warrant further exploration, such as the efficacy of multi-tier tests in quantum physics and developing a cohesive quantum curriculum.
-This paper aims to inform educators and curriculum developers, offering practical recommendations and setting a research agenda to improve conceptual understanding in classical and quantum physics.
+Students often begin physics courses with misconceptions rooted in everyday experience and intuition, which can be resistant to change.
+While research has identified strategies for addressing misconceptions across physics, it remains unclear whether different domains, like classical and quantum physics, require different approaches.
+On one hand, quantum concepts are further removed from daily experience, possibly requiring specialized strategies.
+On the other, all physics must be learned rather than innately understood, and classical physics already contains many counterintuitive ideas, suggesting the same strategies may be equally valid.
+To explore this question, we first develop a structure to organize the existing literature on addressing misconceptions in physics education.
+We identify 126 studies, which we group into six categories, three of which include further subcategories.
+Rather than offering a comprehensive literature review, this scheme provides a structural lens through which to compare how various strategies align.
+We then take an instructor-centered approach, using our framework to guide interviews with quantum physics instructors.
+We interview 12 instructors from the University of Waterloo’s Institute for Quantum Computing and the Perimeter Institute who have collectively taught over 100 quantum courses.
+Our findings suggest that quantum physics instructors find strategies similar to those used in classical physics effective for identifying and addressing misconceptions.
+This highlights the potential to adapt, rather than replace, existing instructional tools to better meet the conceptual challenges of quantum physics.
+Our work provides practical insights for educators and curriculum designers aiming to support deeper understanding in quantum mechanics.
 I Introduction
-Physics is filled with concepts that challenge common intuition and are difficult for students to grasp.
-For example, many students mistakenly believe that heavier objects fall faster than lighter ones or that a constant force is needed to keep an object in motion.
-Such misconceptions often arise from everyday experiences and intuitive reasoning, leading to persistent barriers in learning even after formal instruction [1, 2].
-Misconceptions can be more detrimental than gaps in knowledge.
-This is because students tend to resist new information that conflicts with their existing beliefs, resulting in cognitive dissonance and frustration [3, 4].
-Therefore, effective teaching strategies that identify and rectify these misconceptions are essential for fostering a conducive learning environment.
-However, physics educators face significant challenges in identifying and rectifying these misconceptions.
-Traditional instructional methods often fail to address the underlying intuitive beliefs that give rise to misconceptions [5, 6].
-Not all traditional teaching methods teach for conceptual change [7].
-Consequently, educators must employ innovative and targeted strategies to help students develop a more accurate conceptual framework.
-Fortunately, a sizeable body of research has been conducted on evidence-based strategies for identifying and addressing misconceptions.
-The first aim of this manuscript is to review the existing literature on misconceptions in university physics, focusing on their sources, diagnosis, and remediation strategies.
-Most physics education is centred on classical physics; naturally, most physics education research also focuses on classical physics.
-This trend extends to the study of misconceptions.
-For instance, Gilbert and Watts [1] reviewed 33 studies from 1972 to 1982 on identifying misconceptions, and Soeharto et al.
-[2] reviewed another 111 articles from 2015 to 2019.
-Of the over 40 misconceptions identified in these studies, none were on quantum physics, highlighting a potential gap in the research.
-Quantum physics introduces a unique set of challenges due to its abstract and non-intuitive nature [8, 9, 10].
-Phenomena such as wave-particle duality and entanglement are far removed from everyday experiences and defy classical logic.
-This uniqueness underscores the need for increased attention to identifying and remedying misconceptions in quantum physics.
-The importance of addressing these misconceptions is further amplified by the growing significance of quantum technologies in fields such as computing, cryptography, and materials science.
-This growth has resulted in learners from a broad range of academic backgrounds looking to learn the basics of quantum physics.
-Strategies designed to diagnose and remedy misconceptions in classical physics often rely on analogies and experiences from the macroscopic world, which do not always translate well to quantum phenomena.
-Consequently, while it is essential to determine which ideas and tools from classical physics education can be adapted for quantum physics, it is equally important to recognize that some may not directly apply.
-To explore the applicability of these strategies, we conducted interviews with quantum physics instructors at the University of Waterloo, a leading institute for quantum physics which houses the Institute for Quantum Computing and has strong affiliations with the Perimeter Institute for Theoretical Physics.
-Collectively, these instructors have taught over 100 quantum physics courses.
-Since they are on the front lines of teaching and observing student misconceptions, they provide valuable insights into the effectiveness of existing educational tools and the unique challenges faced in teaching quantum concepts.
-These interviews aim to identify which educational strategies can be ported over from classical to quantum physics.
-This is the second aim of this manuscript.
-The rest of the manuscript is organized as follows.
-Section II presents our review of the literature on addressing misconceptions in university physics.
-Instead of providing an exhaustive survey, we highlight key results and illustrative examples of each.
-A summary of the interviews with quantum physics instructors is presented in Section III.
-By integrating insights from these interviews with the existing literature on classical physics misconceptions, we aim to provide an overview of best practices for addressing misconceptions in classical and quantum physics and identify research opportunities.
-This integrated discussion is presented in Section IV.
-II Review
-This section reviews the literature on misconceptions in university physics education, beginning with a review of the broader context of potential sources of misunderstanding (Section II.1).
-In Section II.2, we explore the origins of these misconceptions.
-Next, we examine methods for diagnosing misconceptions in Section II.3.
-Finally, in Section II.4, we discuss strategies for remedying these misconceptions.
-We propose that these remediation strategies can be categorized into three main types: experiment-based, discussion-based, and reflection-based.
-This categorization helps to identify common elements and key factors contributing to their effectiveness.
-II.1 Misconceptions among other sources of misunderstanding
-The study of misconceptions begins with defining them precisely and distinguishing them from other sources of misunderstanding.
+Physics education research has long grappled with the challenge of student misconceptions (also known as “alternative conceptions,” “conceptual misunderstandings,” or “preconceptions”).
+These are deeply held, often intuitive beliefs that conflict with formal physics understanding.
+Rather than random errors, they represent robust mental models grounded in everyday experience and persist despite formal instruction.
+Over the years, researchers have identified many of these common misconceptions and developed evidence-based strategies to confront and revise them [1, 2, 3].
+An open question is whether classical and quantum physics courses demand distinct strategies for identifying and addressing misconceptions.
+On one hand, quantum concepts are more abstract and less tied to everyday experience, potentially requiring novel instructional approaches.
+On the other hand, all physics—classical or otherwise—must be learned rather than innately grasped, and classical physics also contains counterintuitive ideas, implying that existing methods may still be effective.
+While substantial work has identified misconceptions in both classical and quantum domains, few studies have systematically compared how these misconceptions emerge and are addressed across fields, leaving instructors without clear guidance on whether quantum physics instruction truly demands unique strategies or merely different emphases.
+This study helps address that gap by investigating how highly experienced quantum physics instructors view the strategies for dealing with misconceptions, and whether these strategies must differ from those commonly used in classical contexts.
+First, we construct a framework to organize the existing literature on misconception-oriented instruction, drawing on 126 studies across various subfields.
+This framework then guides a qualitative investigation of expert teaching practice: we conduct structured interviews with 12 quantum physics instructors at the University of Waterloo’s Institute for Quantum Computing and the Perimeter Institute, who collectively have taught over 100 courses.
+Our research is guided by three main questions:
+  • RQ1:
+    What sources and strategies have been identified in the physics education literature for diagnosing and addressing student misconceptions, and how can they be systematically categorized?
+  • RQ2:
+    Do instructors find these strategies effective in quantum physics courses?
+  • RQ3:
+    How do the sources and nature of misconceptions in quantum physics compare to those in classical physics from the perspective of instructors?
+By addressing these questions, we aim to clarify whether teaching quantum physics requires fundamentally new pedagogical methods, or whether existing approaches can be adapted to meet its unique conceptual demands.
+To situate our study, we first highlight key contributions in the literature on student misconceptions in quantum physics instruction.
+A foundational review by Singh and Marshman (2015)[4] synthesized research on common difficulties in upper-level quantum mechanics, emphasizing prevalent reasoning errors, conceptual overgeneralizations, and the importance of research-driven instructional strategies.
+Complementing this, a 2017 review by Krijtenburg-Lewerissa et al.[5] examined teaching practices at the secondary and introductory undergraduate levels, identifying shared misconceptions, gaps in pedagogical approaches, and a lack of validated assessment tools.
+Since these reviews, the field has seen a marked increase in targeted studies.
+Several have focused on specific conceptual challenges students face, such as difficulties with separation of variables [6], solving bound and scattering state problems [7], and applying delta functions in quantum mechanics [8].
+Parallel work has explored strategies for conceptual remediation across a range of foundational topics, including single-photon interference in Mach-Zehnder interferometers [9], the formalism and postulates of quantum mechanics [10], the interpretation of the double-slit experiment [11], and degenerate perturbation theory in the context of the Zeeman effect [12].
+These efforts have contributed to the development of research-based tools such as inquiry-driven tutorials, interactive simulations, and conceptual surveys designed to enhance learning and measure instructional effectiveness.
+In addition, some studies have adopted epistemological frameworks to analyze how students frame and engage with quantum mechanics problems [13, 14].
+Our study expands on this body of research in two key ways.
+First, we center our analysis on the perspectives of experienced instructors, offering a complementary viewpoint.
+Second, we compare strategies across classical and quantum domains to clarify the extent to which established strategies remain effective when addressing quantum physics.
+The rest of this paper is organized as follows: Section II explains the methodology for identifying relevant literature and conducting instructor interviews.
+In Sec.III, we clarify the distinction between misconceptions and other sources of misunderstanding.
+SectionIV details our classification scheme for categorizing the sources of misconceptions, diagnosis methods, and remediation strategies, with illustrative examples.
+Section V then presents the interview findings, highlighting instructors’ insights into the nature of quantum misconceptions and the perceived efficacy of various instructional strategies.
+Finally, Section VI synthesizes these findings to evaluate the transferability of classical teaching approaches to quantum contexts and outlines directions for future research.
+II Method
+II.1 Literature Search
+To conduct a systematic review of misconceptions in university physics education, we employed a multi-stage literature search.
+Our approach combined database searches, structured screening, and citation-based expansion.
+First, we search the Web of Science database using the exact string:
+ALL = ("misconception" OR "alternative conception" OR "preconception" OR "conceptual misunderstanding") AND ALL = ("physics")
+This generated $419$ results and was carried out in January 2024. The first exclusion criteria was
+  • The article had not been published in a peer-reviewed journal.
+This reduced the list to $219$ entries. Each of these $219$ was then independently screened based on the following exclusion criteria.
+  • The article is not in the field of physics education research.
+  • The article is unrelated to addressing misconceptions.
+These criteria reduced the list to $102$ entries. To identify overlooked studies, we employed forward and backward citation tracking:
+  • Backward snowballing: Reviewing references of selected studies.
+  • Forward snowballing: Using Google Scholar’s citation function to find newer papers citing the selected studies.
+This added $24$ papers, bringing the total to $126$.
+Each of these $126$ papers was then sorted into one of four categories: Sources of misconceptions, Diagnostic methods, Remediation strategies, and Other.
+The categories were further subdivided into subcategories (see Fig.
+1).
+Each paper, with each subcategory and a brief summary is provided in Appendix B.
+Figure 1: Overview of our organizational scheme for the literature on misconceptions in physics.
+Numbers indicate the count of papers in each category.
+For brief summaries of each paper, see Tables 2 to 7 in Appendix B.
+II.2 Interviews
+II.2.1 Participant Selection
+We focused on the University of Waterloo due to its recognized strengths in quantum physics education and its affiliation with the Perimeter Institute.
+The university offers 16 undergraduate and 37 graduate courses on solely quantum physics and 9 undergraduate courses covering quantum physics (listed in Appendix A).
+Initial participants were identified through official course listings; they were contacted by email using a standardized invitation.
+We also used snowball sampling by asking interviewees to recommend colleagues with relevant experience.
+In total, 12 instructors (nine current faculty members and three experienced guest lecturers for summer schools) agreed to participate.
+Each participant had taught quantum physics in a university setting for a minimum of five years, ensuring substantial teaching experience.
+II.2.2 Interview Protocol
+We designed our interview protocol following guidelines from qualitative research in education [15, 16, 17, 18]. The final protocol included five core, open-ended questions addressing:
+  • What are university students’ common misconceptions about quantum physics?
+  • What are the sources of these misconceptions?
+  • What are your strategies for identifying misconceptions?
+  • What are your strategies for addressing misconceptions?
+  • Besides misconceptions and gaps in knowledge, what else causes students to make mistakes?
+We asked follow-up questions as relevant, while ensuring all core questions were covered to maintain consistency across interviews.
+Each interview was scheduled for 30 minutes via Zoom, though many extended up to 45 minutes or more based on participant engagement.
+Interviews were audio-recorded and professionally transcribed.
+All participants provided recorded verbal consent.
+We assured them of confidentiality; hence, all identifying information was removed from transcripts.
+II.2.3 Strengths and Limitations of the Interview Approach
+Interviews with lecturers are not the standard method for identifying students’ difficulties.
+As a non-standard approach, we carefully examine its strengths and weaknesses.
+One strength is that lecturers, through years of teaching, develop an intuitive understanding of common student struggles, providing a broad overview of recurring challenges.
+Additionally, unlike standardized tests or surveys, interviews allow lecturers to reflect on not only errors in assessment but also patterns in classroom engagement, misconceptions, and areas where students frequently seek help.
+Moreover, interviews can be more time-efficient than large-scale student surveys or detailed diagnostic assessments.
+However, this method has notable weaknesses.
+Lecturer recollections of student difficulties may be influenced by selective memory or biases.
+The identification of challenges is subjective and may vary between lecturers based on their personal experiences, teaching styles, and interactions with students.
+Furthermore, some student difficulties may not be apparent to lecturers, particularly if students do not actively express them in class.
+Lastly, because this method does not directly involve students, it risks overlooking nuanced challenges that students themselves perceive but lecturers do not recognize.
+To ensure the validity of our findings with this strategy, we implemented several procedures.
+First, we triangulated interview data with a literature review to strengthen our conclusions.
+Second, we standardized the interview process by using structured questions and a systematic approach, focusing on specific topics and common misconceptions to reduce arbitrariness.
+Third, we gathered insights from multiple lecturers teaching the same subject to identify consistent patterns and enhance the robustness of our findings.
+These measures help mitigate the limitations of lecturer interviews and improve the reliability of our conclusions.
+III Defining “Misconception”
+Before presenting our main results, we define what is meant by a “misconception” by comparing it against other sources of misunderstanding.
 In their 1983 review, Gilbert and Watts noted the initial lack of consensus on the definition of a misconception within the field [1].
@@
 For example, translating a word incorrectly is a vernacular misunderstanding that can likely be corrected easily without leading to deeply ingrained incorrect beliefs.
-Having clearly defined misconceptions and tied our definition to the existing ones used in the literature, we turn to their sources.
-II.2 Sources of misconceptions
-Misconceptions can arise from three sources:
+IV Categorizing the literature
+This section addresses RQ1, by creating broad categories to organize the literature based on the papers we found via our literature search.
+This categorization helps to identify common elements and key factors contributing to their effectiveness.
+In Section IV.1, we explore the origins of these misconceptions.
+Next, we examine methods for diagnosing misconceptions in Section IV.2.
+Finally, in Section IV.3, we discuss strategies for remedying these misconceptions.
+For each subcategory we present at least one illustrative example.
+IV.1 Sources of misconceptions
+The literature we found on sources of misconceptions is presented and sorted in Table 2. Misconceptions can arise from three sources:
   • Sources of knowledge: Misconceptions can arise from inaccurate information from sources students believe they should trust to some extent.
@@
 This is potentially due to the abstract nature of quantum concepts exacerbating the impact of vague instructional materials.
-II.3 Diagnostic Methods
+IV.2 Diagnostic methods
 Being mindful of sources of misconceptions will reduce the number of students’ misconceptions, but some will emerge.
 To help students address these misconceptions, we need tools to diagnose them.
+The literature we identified on diagnostic methods is presented and sorted in Table 3.
 The commonly used tools are interviews, open-ended questions, certainty of response index (CRI) tests, and multiple-tier tests.
 The first two strategies are typically used in courses with few students, and the latter two in courses with many students.
-Interviews and open-ended questions
+IV.2.1 Interviews and open-ended questions
 Interviews involve one-on-one interactions between an interviewer and a student.
@@
 Additionally, students may not provide detailed answers unless incentivized, which can limit the effectiveness of this method compared to oral interviews.
-CRI and multiple-tier tests
+IV.2.2 CRI and multiple-tier tests
 The foundational concept of "multiple-tier" physics tests was proposed by Hasan et al.
@@
 Multi-tier tests may be more effective, or physics education research may have drifted towards studying classrooms with more students.
-II.4 Remediation Strategies
+IV.3 Remediation strategies
 After a misconception is identified, it can be addressed.
@@
 Examples for each category are summarized in Table 1.
+All of the literature we examined in this category is presented in Table 4 and 5.
 Strategy Explanation
@@
 Table 1: Summary of educational strategies categorized into reflection-based, discussion-based, and experiment-based approaches.
-Reflection-based strategies
+IV.3.1 Reflection-based strategies
 Reflection-based strategies encourage students to introspect and critically reassess their knowledge, integrating new and accurate information.
@@
 Their results suggest that the act of making predictions enhances students’ observational accuracy and helps them retain the correct outcomes.
-Discussion-based strategies
+IV.3.2 Discussion-based strategies
 Discussion-based activities use social interaction to boost learning.
@@
 Similarly, following peer discussions, correct responses for the net work done in a cyclic process rose from 28% to 51%, and for heat in a cyclic process, from 15% to 40%.
-Experiment-based strategies
+IV.3.3 Experiment-based strategies
 The final category, experiment-based activities, emphasizes “hands-on” engagement, allowing students to interact directly with physical or simulated systems.
@@
 Virtual and physical sequences improved understanding, but virtual interactions alone were more impactful than physical ones alone, emphasizing the potential of simulation-based learning tools in educational settings.
-III Experiences from quantum physics educators
+V Interviews with quantum physics educators
 Most of the tools for identifying and addressing misconceptions were developed for classical physics.
-Since quantum physics education presents unique challenges, we need to consider the applicability of these tools to this subject.
+Since quantum physics education may present unique challenges, we need to consider the applicability of these tools to this subject.
 We conduct such an evaluation through long-form interviews with quantum physics instructors at the University of Waterloo.
-This study aims to identify misconceptions in quantum physics and assess effective tools for diagnosing and rectifying these misconceptions.
-III.1 Methodology Overview
-First, we outline our methodology for identifying instructors and conducting interviews.
-The University of Waterloo, renowned for its quantum physics research and extensive course offerings, was chosen for this study.
-It houses the Institute for Quantum Computing and is closely affiliated with the nearby Perimeter Institute for Theoretical Physics.
-A testament to the university’s focus on quantum physics is the volume of courses it offers on the subject.
-Listed in Appendix A, the university offers 16 undergraduate and 37 graduate courses on solely quantum physics and 9 undergraduate courses covering quantum physics.
-We identified faculty members who taught the 16 undergraduate courses via online course listings and departmental contacts.
-Emails were sent to these faculty members requesting interviews, resulting in nine agreements.
-Through these nine, another faculty member was identified to give an interview about their experience teaching one of the nine undergraduate courses partially covering quantum physics.
-Additional interviews were secured with two individuals from the Institute for Quantum Computing and Perimeter Institute with years of experience teaching quantum physics to undergraduate students.
-The twelve instructors have collectively taught over 100 university quantum physics courses and many summer schools, reaching thousands of students.
-With no detailed guide for conducting physics education research interviews, we relied on guides from other fields [43, 44, 45, 46].
-The interviews aimed to identify misconceptions in quantum physics and effective tools for addressing them.
-Five core questions were developed, with flexibility for follow-up questions based on responses.
-Verbal consent for recording was obtained, and confidentiality and withdrawal rights were explained to each interviewee.
-Interviews were scheduled for 30 minutes each, conducted via Zoom, and many were extended beyond the scheduled time.
-The insights gathered from these interviews are synthesized and presented in a cohesive narrative in the following subsection.
-III.2 Synthesis of Insights
-We present each question and list the various answers instructors gave. With few exceptions, each answer is an amalgamation of comments given by multiple instructors.
-What are university students’ common misconceptions about quantum physics?
+These interviews aim to address RQ2 and RQ3.
+We present each question and list the various answers instructors gave.
+With few exceptions, each answer is an amalgamation of comments given by multiple instructors.
+V.1 What are university students’ common misconceptions about quantum physics?
 Entanglement.—Students often believe that entanglement implies faster-than-light communication.
@@
 Misinterpreting photon interactions and energy levels can lead to significant misunderstandings in quantum optics exercises.
-What are the sources of these misconceptions?
+V.2 What are the sources of these misconceptions?
 The instructors collectively identified eight sources of misconceptions. The first four pertain to teaching methods, the next three to educational materials, and the last to the students.
@@
 Teaching the limitations of different problem-solving tools and methods can help students apply these tools appropriately in various contexts.
-What are your strategies for identifying misconceptions?
+V.3 What are your strategies for identifying misconceptions?
 Identifying trends in formative assessment errors.—Instructors commonly analyze student responses to exercises to detect misconceptions.
@@
 This consistency allows the instructor to anticipate and proactively address common misunderstandings in future courses.
-What are your strategies for addressing misconceptions?
+V.4 What are your strategies for addressing misconceptions?
 In-class quizzes with “misconception traps”.—Several instructors utilize quizzes designed to address misconceptions.
@@
 However, most instructors pointed out the limited options for this in quantum physics classrooms.
-Besides misconceptions and gaps in knowledge, what else causes students to make mistakes?
+V.5 Besides misconceptions and gaps in knowledge, what else causes students to make mistakes?
 Three instructors said students make mistakes for reasons outside of misconceptions and gaps in knowledge. All three pointed to deficiencies in prior knowledge.
@@
 The scheduling of these courses sometimes fails to account for the necessary preparatory coursework, leading to students encountering complex quantum topics without a solid foundational understanding.
-Other miscellanenous comments
+V.6 Other miscellanenous comments
 Several insightful side conversations emerged from the interviews, providing valuable perspectives on addressing misconceptions.
@@
 This contrasts with the challenge of teaching students with no prior conception.
-Specialization in quantum physics:.—The importance of tailoring quantum mechanics education to the audience and its goals was emphasized.
+Specialization in quantum physics.—The importance of tailoring quantum mechanics education to the audience and its goals was emphasized.
 Mathematics students may approach quantum mechanics primarily through the lens of linear algebra with additional rules.
@@
 However, the question is then raised on what to include in undergraduate curriculums as the number of quantum physics education courses expands.
-III.3 Limitations of the study
-In this work, we focused on one institute that is a hot spot for quantum physics education.
-This intense concentration of experiences made Waterloo a prime location for analyzing quantum physics education.
-However, Waterloo is one of Canada’s most competitive physics programs and is recognized as a global leader in quantum physics.
-As such, the student population is likely biased to have a stronger physics background than the average student.
-Furthermore, the instructors likely understand the content more deeply than those at a randomly chosen university.
-Future research should aim to include a larger and more diverse sample of instructors from various institutions and cultural backgrounds to enhance the generalizability of the findings.
-The study relies heavily on qualitative data from interviews, which, while rich in detail, may introduce subjectivity and bias.
-The interpretations and conclusions drawn are influenced by the educators’ and interviewer’s perspectives and experiences.
-Including quantitative measures, such as surveys or pre- and post-assessments of student performance, could provide a more balanced view and validate the effectiveness of the suggested strategies.
-IV Discussion
+VI Discussion
 Physics educators often face challenges due to ingrained student misconceptions.
-Classical and quantum physics present unique hurdles, as highlighted by the literature and reinforced by our interviews with quantum physics educators at the University of Waterloo.
-This final section synthesizes many of the conceptual insights from our review and discussions and identifies key research questions.
-However, one should refer to the main text for a detailed list of strategies.
-IV.1 Integration of insights from literature and interviews
+Whether classical and quantum physics requires unique strategies for identifying and addressing them is an open question.
+From the interviews we conducted, instructors seem to find the same strategies effective for quantum physics instruction as instruction of other courses.
+We conclude here with a brief summary of the insights this work gleaned from the literature search and interview and identifying open research questions.
+VI.1 Integration of insights from literature and interviews
 Instructors can save considerable time by familiarizing themselves with the common misconceptions in their courses that have already been identified in prior studies.
@@
 We note that experiment-based strategies are less used when teaching quantum physics.
-However, this is due to their lack of availability, not lack of efficacy.
-IV.2 Key research questions
-Multi-tier tests have become prevalent for identifying misconceptions in physics education but are less commonly used in quantum physics education.
+However, this is likely due to their lack of availability, since there is nothing indicating a lack of efficacy.
+VI.2 Key research questions
+Multi-tier tests have become prevalent for identifying misconceptions in physics education but have not been systematically studied in quantum physics education.
 Investigating their effectiveness in this field is a promising research opportunity.
@@
 The quantum physics misconceptions identified by educators at Waterloo are largely conceptual, contrasting with the primarily mathematical and technical misconceptions identified by Styer in 1996 [51].
-A systematic study comparing the relative importance of addressing these different types of misconceptions may also prove helpful in identifying effective remediation strategies.
+A systematic study with groups of student comparing the relative importance of addressing these different types of misconceptions may also prove helpful in identifying effective remediation strategies.
 Instructors highlighted that deficiencies in prior knowledge, particularly in mathematical training and foundational quantum physics, significantly impact student performance.
@@
 Acknowledgements.
-This article is part of my work to fulfill the requirements for the University of Waterloo’s Centre for Teaching Excellence’s Certificate in University Teaching.
-This work is distinct from my PhD research [49, 50, 51, 52, 53, 54, 55, 56, 57, 58] which was funded by the Vanier CGS.
+This work is distinct from my PhD research that was funded by the Vanier CGS [53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
 However, as this article was written during this time, I would like to acknowledge and express gratitude for the support provided by the Vanier CGS.
@@
     QIC 890 Topics in Quantum Information
+Appendix B Database of works
+Ref. Subcategory Summary
+[63] Sources of Knowledge Examines definitional problems in physics education rather than misconceptions, proposing a framework to improve definitions.
+[64] Sources of Knowledge Investigates how different approaches to drawing free body diagrams affect students’ understanding of Newton’s laws.
+[65] Sources of Knowledge Categorizes common misconceptions about sound propagation and acoustic devices.
+[66] Sources of Knowledge Conducts a literature review to classify and list common misconceptions about solid friction.
+[67] Sources of Knowledge Analyzes how textbook representations of optics contribute to student misconceptions.
+[68] Sources of Knowledge Identifies misconceptions about semiconductor phenomena and explores how simulations contribute to their formation.
+[69] Students Examines how language confusion can be mistaken for misconceptions, showing that question misinterpretation can lead to false diagnoses.
+[70] Students Studies linguistic misconceptions rather than physics-related misconceptions, examining how language structure interferes with English sentence translation.
+[71] Students Investigates how students with misconceptions focus on misleading information during physics problem-solving.
+[72] Students Investigates how inhibitory control affects the ability to overcome misconceptions about position-velocity relationships.
+[23] Teaching Methods Examines how visual representations impact cognitive load and learning, emphasizing the role of prior knowledge.
+[73] Teaching Methods Discusses conceptual change in physics without focusing specifically on misconceptions, critiquing traditional expert-novice dichotomies.
+[74] Teaching Methods Identifies persistent Aristotelian misconceptions about motion and force in secondary students.
+[75] Teaching Methods Investigates students’ reluctance to use mental simulations in problem-solving due to fear of misconceptions, rather than identifying specific misconceptions.
+[76] Teaching Methods Argues that many misconceptions across physics topics stem from a broader misunderstanding of probability concepts.
+[77] Teaching Methods Examines cognitive limitations (working memory, inhibition, set-shifting) as sources of misconceptions in physics.
+[78] Teaching Methods Identifies systematic errors in understanding projectile motion, suggesting an intuitive impetus theory as the cause.
+Table 2: Summary of papers that were categorized under “Sources of misconceptions”.
+Ref. Subcategory Summary
+[79] CRI and MTT Develops a diagnostic test to assess students’ misconceptions in quantum physics.
+[80] CRI and MTT Develops and applies a three-tier test to assess misconceptions about heat and temperature across different education levels.
+[81] CRI and MTT Develops a three-tier test to diagnose misconceptions about velocity and force in circular motion.
+[82] CRI and MTT Develops a concept inventory to assess misconceptions about heat, energy, and thermal radiation, showing misconceptions are highly persistent.
+[83] CRI and MTT Develops and analyzes a relativity concept inventory, examining misconceptions and confidence in answers.
+[84] CRI and MTT Develops and validates a three-tier diagnostic test for misconceptions about heat, temperature, and internal energy.
+[85] CRI and MTT Develops and implements a three-tier diagnostic test to assess misconceptions about the photoelectric effect.
+[86] CRI and MTT Develops and validates a four-tier test to identify misconceptions in geometrical optics.
+[87] CRI and MTT Uses a four-tier diagnostic test to assess misconceptions about liquid pressure, finding that misconceptions persist even among future teachers.
+[88] CRI and MTT Develops and validates a three-tier test to assess misconceptions about work, power, and energy, revealing common alternative conceptions.
+[89] CRI and MTT Uses Rasch measurement to analyze misconceptions across multiple science disciplines and identify item difficulty patterns.
+[27] CRI and MTT Identifies student misconceptions in mechanics using a three-tier diagnostic test and clinical interviews.
+[26] CRI and MTT Develops and validates a four-tier diagnostic test to identify student misconceptions in fluid mechanics.
+[25] CRI and MTT Develops and validates a four-tier diagnostic test with confidence measures to identify student misconceptions in fluid mechanics.
+[28] CRI and MTT "Investigates prospective teachers’ confidence-accuracy calibration in diagnosing potential difference misconceptions. "
+[90] I and OEQ Develops and administers two-tier tests to diagnose misconceptions in mechanics, electricity, heat, waves, and optics, revealing persistent misconceptions across age groups.
+[91] I and OEQ Uses multi-dimensional cognitive analysis to categorize students’ understanding of heat conduction and identify naive ideas.
+[92] I and OEQ Develops and validates concept questions to diagnose students’ misconceptions about galaxies and spectra.
+[93] I and OEQ Uses visual representations and interviews to diagnose students’ preconceptions about friction.
+[94] I and OEQ Focuses on methods for analyzing student misconceptions using TIMSS data.
+[95] I and OEQ Develops a system for diagnosing students’ cognitive states and preconceptions to detect misconceptions.
+[96] Other: Alluvial diagrams Uses alluvial diagrams and subquestions to analyze depth of conceptual understanding in Newtonian mechanics.
+[97] Other: Cluster analysis Uses cluster analysis to identify patterns in students’ misconceptions about force and Newton’s third law.
+[98] Other: Module analysis Uses module analysis to identify patterns in misconceptions about electricity and magnetism.
+[99] Other: Module analysis Uses module analysis to identify and compare persistent misconceptions in Newtonian mechanics across different institutions.
+[100] Other: New tool Develops an instrument to assess students’ understanding of the relationships between force, velocity, and acceleration.
+Table 3: Summary of papers that were categorized under “Diagnostic methods”. “I and OEQ” is interviews and open-eded equations. “CRI and MTT” is CRI and multiple-tier tests.
+Ref. Subcategory Summary
+[101] Discussion Evaluates conceptual assignments and discussions as methods to improve understanding and reduce misconceptions about force and motion.
+[102] Discussion Develops and evaluates a technology-enhanced learning environment to address misconceptions about seasons.
+[103] Discussion Uses a bridging analogy-based instruction strategy to address misconceptions about Newton’s Third Law while considering gender and group effects.
+[104] Discussion Uses constructivist worksheets to successfully address misconceptions about force and motion.
+[105] Discussion Compares problem-based learning with traditional methods, finding that problem-based learning is more effective at reducing misconceptions.
+[106] Discussion Examines the effectiveness of the 7E model in addressing misconceptions in electricity compared to traditional methods.
+[107] Discussion Compares traditional instruction with the 4MAT teaching method, showing it is more effective in reducing misconceptions about electricity.
+[108] Discussion Evaluates the effectiveness of concept cartoon-embedded worksheets in correcting misconceptions about Newton’s Laws.
+[109] Discussion Evaluates the use of concept cartoons in addressing alternative conceptions about chemical bonding within a context-based approach.
+[39] Discussion Uses hints and peer discussion in a lecture-based intervention to help students overcome misconceptions in thermal physics.
+[41] Discussion Uses multimedia to highlight and correct misconceptions in physics, helping students recognize inconsistencies in their reasoning.
+[40] Discussion Examines how online collaborative learning affects middle school students’ science misconceptions, finding increased misconceptions.
+[110] Experiment Compares lab-based and computer-assisted instruction, finding computer-assisted education more effective in addressing misconceptions about force and motion.
+[111] Experiment Develops interactive multimedia to address misconceptions about dynamic fluids through cognitive conflict-based learning.
+[43] Experiment Develops and evaluates a virtual physics system that customizes tutoring to address misconceptions in mechanics and energy concepts.
+[48] Experiment Uses virtual 3D solar system simulations on tablets to improve students’ understanding of astronomical scale and correct misconceptions.
+[46] Experiment Uses video analysis and interactive software to correct student misconceptions about car braking distances.
+[44] Experiment Uses dynamic simulations to reduce misconceptions about CO2 stabilization, showing improved understanding through experience-based learning.
+[47] Experiment Compares physical and virtual simulations in teaching electric circuits, showing both reduce misconceptions but with persistent difficulties.
+[49] Experiment Compares immersive and non-immersive virtual environments for teaching relative motion, finding that immersion enhances learning of 2D motion.
+[42] Experiment Compares traditional, mental modeling, and analogy-based teaching methods for improving student understanding of electric circuits.
+[45] Experiment Examines the effectiveness of simulations for teaching electricity concepts to young students, highlighting benefits and potential misconceptions.
+[112] Experiment Compares the effectiveness of cognitive perturbation vs. cognitive conflict in facilitating conceptual change in electricity and magnetism.
+[113] Experiment Examines the effectiveness of the ICI model with simulations in addressing misconceptions about force and vibration.
+[114] Other: Instructional Explores a field model as an alternative to traditional electron flow models, showing improved understanding of direct current.
+Table 4: Summary of papers that were categorized under “Remediation strategies”. The table was too large to fit one one page, so the second half is in the subsequent table.
+Ref. Subcategory Summary
+[115] Reflection Compares conceptual change methods (texts, analogies, computer animations) and finds that a combination approach is most effective.
+[116] Reflection Identifies misconceptions about color and tests concept mapping as a strategy to address them, showing long-term learning benefits.
+[117] Reflection Uses concept mapping to reveal misconceptions in geometric optics among first-year university students.
+[118] Reflection Explores how misconceptions about gravitational acceleration can be leveraged to enhance understanding.
+[119] Reflection Evaluates conceptual change-oriented instruction using conceptual change texts and cartoons, showing improved understanding but no attitude change.
+[120] Reflection Addresses a misconception in relativity and proposes using the twin paradox as a corrective instructional strategy.
+[121] Reflection Investigates how analogies and examples can be used effectively to correct misconceptions, emphasizing explicit analogy development.
+[36] Reflection Develops and introduces a computer-supported conceptual change text to address misconceptions in fluid pressure.
+[35] Reflection Compares conceptual change instruction with traditional methods for improving students’ understanding of heat and temperature.
+[37] Reflection Evaluates the effectiveness of conceptual change texts and experiments in addressing misconceptions about matter and change.
+[33] Reflection Examines how model-based teaching affects pre-service teachers’ mental models of lunar phenomena.
+[38] Reflection Investigates how epistemic beliefs and knowledge representations affect cognitive processing and conceptual change in physics.
+[34] Reflection Analyzes how students’ prior knowledge influences their perception and retention of physics lecture demonstrations.
+[32] Reflection Explores a writing-based strategy (folder activity) to help students identify and address physics misconceptions.
+[122] Reflection Studies how refutational texts can induce conceptual change, helping to correct misconceptions.
+[123] Reflection Extends the concept of refutation texts beyond science education, showing benefits in history learning.
+[124] Reflection Examines the effectiveness of explicit vs. implied refutations in science texts, showing that even implied refutations help reduce misconceptions.
+[125] Reflection Compares different online learning tools and their effectiveness in reducing misconceptions about free-fall motion.
+[126] Reflection Explores how refutational text structure influences students’ conceptual change regarding counterintuitive physics concepts.
+Table 5: Summary of papers that were categorized under “Remediation strategies”. The table was too large to fit one one page, so the first half is in the previous table.
+Ref. Summary
+[127] Identifies a misconception that nuclear attraction is evenly shared among electrons, highlighting issues at the physics-chemistry boundary.
+[128] Identifies misconceptions about inertia, gravity, and gravitational force among pre-service teachers.
+[129] Demonstrates experimental confirmation of a common misconception in the application of Einstein’s photoelectric equation.
+[130] Identifies 32 misconceptions about radiation and radioactivity, comparing students from different disciplines.
+[131] Identifies 38 misconceptions about force and 15 about acceleration, along with four root causes.
+[132] Identifies alternative conceptions about kinematics, particularly related to the meaning of zero in displacement, speed, velocity, and acceleration.
+[133] Identifies persistent alternative worldviews in students’ Force Concept Inventory responses before and after instruction.
+[134] Uses international assessment data to identify misconceptions, errors, and misunderstandings in physics and mathematics.
+[135] Identifies misconceptions about the causes of seasons, with a focus on inconsistencies in their reasoning.
+[136] Examines cognitive resistance to correcting misconceptions about air pressure and vacuums, showing how learners modify knowledge to maintain misconceptions.
+[117] Identifies misconceptions about electric current, including potential difference, generators, and magnetic fields, with gender differences in understanding.
+Table 6: Summary of papers that were categorized under “Examples of Misconceptions”.
+Ref. Summary
+[137] Documents the persistence of the misconception that vision involves emissions from the eyes, a belief rooted in historical theories.
+[138] Explores how misapplications of Laplace’s law lead to misunderstandings about alveolar inflation in physiology education.
+[139] Uses the Force Concept Inventory (FCI) to diagnose misconceptions about force and motion, finding strong misconceptions about impetus and active force.
+[140] Examines why misconceptions about emergent processes (e.g., diffusion, electricity, heat, and evolution) are more robust than misconceptions about direct processes.
+[141] Discusses difficulties in interpreting quantum mechanics due to mixing classical concepts with quantum ones, leading to misconceptions.
+[142] Addresses a common misconception about thermal behavior of materials, explaining the difference between thin and thick slabs when heated.
+[143] Uses analogy-based testing to analyze misconceptions in Newton’s Third Law, finding gender-based and class-level differences.
+[144] Examines how different task formats influence understanding of accelerated motion, showing a shift from linear to accelerated motion understanding.
+[145] Investigates misconceptions about the microscopic and wave nature of sound across different engineering disciplines.
+[146] Investigates intuitive misconceptions about velocity changes, focusing on the "changes-take-time" misconception and its causes.
+[147] Categorizes students’ alternative conceptions based on epistemological and ontological frameworks.
+[148] Compares scientific reasoning and knowledge between those with persistent misconceptions and those with accurate conceptions.
+[149] Finds that future teachers struggle to differentiate between attribute and measurement, impacting their ability to teach these concepts effectively.
+[150] Tracks how children’s understanding of projectile motion evolves, showing early influences of contextual features and later adoption of impetus misconceptions.
+[151] Investigates why students avoid answering conceptual questions and whether learning occurs in a response system setting.
+[152] Analyzes incorrect answers in the Force Concept Inventory to understand robust non-Newtonian worldviews.
+[153] Tracks the development and persistence of thermal conceptions across grade levels, showing some misconceptions persist despite education.
+[154] Examines the persistence of materialistic thinking in sound concepts and how misconceptions evolve across academic levels.
+[155] Compares misconceptions about impetus theory and projectile motion between students with and without autism, showing differences in adoption rates.
+[156] Compares misconceptions about electric circuits across different education levels.
+[157] Investigates how explainer videos with misconceptions can create an illusion of understanding, leading to persistent misconceptions.
+[158] Examines whether confidence in incorrect answers correlates with misconceptions about radioactive decay.
+[159] Diagnoses misconceptions about force and motion, showing differences based on sex and grade level, with implications for targeted remediation.
+[160] Examines misconceptions in scientific reasoning using confidence-based assessment, showing high confidence in incorrect explanations.
+[161] Differentiates between misconceptions and a mere lack of knowledge in understanding electric circuits.
+[162] Investigates misconceptions about AI trustworthiness in physics learning, showing a high level of misplaced trust.
+[163] Examines misconceptions in astronomy and how they relate to interest in the subject.
+[164] Investigates how schooling differences impact geometric misconceptions, finding that ultraorthodox students sometimes outperform their mainstream peers.
+Table 7: Summary of papers that were categorized under “Nature of misconceptions”.